OpenStudy (sanchez9457):

Find the particular solutions of the following ODE: y''+2y'-3y=e^t

OpenStudy (sanchez9457):

\[y''+2y'-3y = e ^{t}\]

OpenStudy (nincompoop):

let us see your solution

OpenStudy (sanchez9457):

\[y = e ^{t}\] ?

OpenStudy (anonymous):

\[ e^t+2e^t-3e^t = 0\neq e^t \]

OpenStudy (anonymous):

You may want to try: \[ Ae^{Bt} \]

OpenStudy (anonymous):

\[ y = Ae^{Bt}\\ y' = BAe^{Bt}\\ y'' = B^2Ae^{Bt}\\ \]

OpenStudy (anonymous):

Actually it's probably just \(y=Ae^t\) given how we know \(B=1\).

OpenStudy (anonymous):

you have to multiply by x since r is positive 1

OpenStudy (anonymous):

I suppose it could be \(Ae^t+Bte^t\)

OpenStudy (anonymous):

well i believe you will have Ae^-3t+Be^t+ z(t) and then you need to solve for the particular solution.

OpenStudy (anonymous):

and when you solve for z(t) you will need to multiply by t, because you already have a e^t in the general solution.

OpenStudy (sanchez9457):

It's actually y=Ate^t. If I had y=e^t I would get 0=0. Thanks though!